# Textoshop: Interactions Inspired by Drawing Software to Facilitate Text Editing

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; copyright: noneccs: Human-centered computing Interaction tech
@@
This is the teaser description for screen readers.
@@ 1. Introduction
The processes involved in writing share many similarities with those involved in image editing. Just as writers plan and refine a story by outlining the narrative and rearranging the story flow (Flower and Hayes, 1977), visual artists plan and refine a graphical composition by laying out visual elements on a canvas, rearranging them, sometimes organizing them in layers  (Tsandilas et al., 2015). Just as writers combine arguments into sentences and paragraphs (Oshima and Hogue, 2005), visual artists combine shapes into complex drawings (Wood, 2022). And just as writers revise the language and continuously experiment with wording and structure (Faigley and Witte, 1981), visual artists revise the aesthetics and continuously experiment with colours and textures (Tsandilas et al., 2015).
@@
To better support writers [^1] By “writers”, we do not necessarily mean professional writers but any individual who engages in writing activities, no matter their expertise, we propose to adapt solutions used in drawing software to help with recurring writing tasks. Specifically, we aim to (1) capitalize on solutions that have been proven to work well and (2) offer solutions most users are already familiar with. Loosely, we view words as pixels, sentences as regions, and tones as colours. This analogy resulted in the design of Textoshop (Figure 1), a text editing software that supports direct manipulation to move, resize and restructure text selections; a layer management system to organize, display, and edit text in independent views; a sentence composition feature to construct sentences or paragraphs by combining, intersecting, or subtracting ideas or phrases; a tone picker to explore different tones when editing text; and tools including an “eraser” to intelligently remove part of the text, a “tone eyedropper” to copy the tone of a passage, and a “smudge” tool to paraphrase and smoothen transitions between sentences.
@@ 2. Background and Related Work
Traditional text editors such as Microsoft Word offer limited support for editing the text. We first review the issues this creates before discussing solutions proposed in the literature. We then detail the rationale behind our approach to take inspiration from drawing software [^2] Throughout this paper, we use the term “drawing software” to refer to both raster (e.g., Adobe Photoshop) and vector (e.g., Inkscape and Adobe Illustrator) graphics editors. .
@@ 2.1. Issues with Traditional Ways of Editing Text
Second, exploring the space of possible text variations is difficult. A sentence can be rewritten in many ways by varying tone, word order, or length. In fact, rewriting is often considered “the essence of writing” (Zinsser, 2016). Yet, most text editors offer little support to help revise language other than a thesaurus to find synonyms. Some experimental tools leverage LLMs to revise sentences or paragraphs, but it remains challenging for writers to express their intent with fine granularity (Kim et al., 2024).
@@ 2.2.1. Writing Using Non-linear Text Editors
As early as the development of the World Wide Web, Bolter and Joyce (1987) proposed a writing tool supporting interactive fiction writing using a structural editor with a diagrammatic view. This spurred research into the development of hypertext authoring tools (Kitromili et al., 2019) with which authors write part of the story in an order that might differ from how readers will read it. While these tools were primarily designed to create interactive stories, the idea of being able to write in any order inspired tools to author more traditional and static text. For example, the commercial software Scrivener (Scrivener, 2005) allows writers to compose their text in any order. Writers can explicitly structure their writing using different text entries and then reorganize these text entries to reorganize the final story.
Textoshop also supports non-linear writing: text can be either positioned following lines on a page or freely arranged in 2D in and off the page. The margins can store mood boards, alternatives, and outlines akin to how digital artists leverage the space outside the canvas to preserve elements for reference (Holinaty et al., 2021). Additionally, Textoshop supports text layers, where elements of text can be edited independently and combined at will to create different composite stories.
@@ 2.2.2. Revising and Exploring Text Variations
Our work differs in that the space of possible variations is explored through direct manipulation. If a sentence is close to satisfactory, users can explore variations of different lengths or different word orders through the resizing and rotating features. They can also move a slider in a two-dimensional space to explore tone nuances, much like visual artists explore colour nuances in a colour picker.
@@ 2.2.3. Reordering and Combining Text Fragments
New writing tools increasingly “shift the writer’s role towards editorial decisions” and manipulating text fragments (Buschek, 2024). For example, Langsmith (Ito et al., 2020) can suggest full sentences from incomplete and rough sentence fragments. Similarly, Rambler (Lin et al., 2024) allows refining and combining dictated train-of-thoughts into full drafts. Other work proposed to manipulate summary representations such as paragraph summaries (Dang et al., 2022) and visual outlines (Zhang et al., 2023) to edit the structure and flow of ideas in a draft.
We go beyond simple merging of text passages by exploring novel ways to combine text inspired by boolean operations on shapes, such as subtracting to remove the information in one passage from another, intersecting to preserve only shared information between two passages, and excluding to remove redundant information redundant between two passages.
@@ 2.3. Designing Interactions Based on Analogies
Designing software interactions based on such analogies has proven successful many times in the past. For example, the direct manipulation interaction paradigm is all about manipulating elements in software as if they were physical objects that can be moved, resized, or distorted (Shneiderman, 1983; Hutchins et al., 1985). Instrumental interaction extends direct manipulation by noting that humans leverage tools when the finger or the hand is not enough (Beaudouin-Lafon, 2000; Beaudouin-Lafon and Mackay, 2000). Translated into the software domain, these tools become interaction instruments that operate on domain objects.
Drawing software combine a canvas analogy with ideas from direct and instrumental interaction. For example, Adobe Photoshop allows elements to be moved, removed, resized, and rotated through rapid and reversible physical actions that mimic how one interacts with physical objects in the real world, just like advocated by direct manipulation. And interaction instruments allow simulating paint brushes, pencils, or erasers. Other interaction instruments include the colour picker, which is analogous to a paint palette, and layers, which are analogous to transparent sheets (Belge et al., 1993; Andrews, 2006) and reify (Beaudouin-Lafon and Mackay, 2000) the concept of independent manipulation.
The success of drawing software has been inspirational for other software. For example, it is common to find the concept of brushes, colour pickers, and layers adopted by software operating in different domains, such as presentation and video editing. And some of the digital drawing concepts have been adapted successfully, such as using a paintbrush to control the fortune of characters in a story (Chung et al., 2022), mixing prompts just like colours can be mixed to convey visual styles difficult to express with words (Chung and Adar, 2023), or repurposing raster image tools to manipulate colour histograms (Chevalier et al., 2012).
In contrast, text editors evolved from command-line word processors (Mahlow, 2023; Haigh, 2006) and use the analogy of the sheet of paper (Kirschenbaum, 2016) and the typewriter. In this work, we explore how interaction inspired from the visual creative domain could open up new possibilities for text editing.
@@ 3. Textoshop
We designed Textoshop by reviewing features found in most raster or vector image editing software and that most computer users are likely familiar with, such as colour pickers, drawing tools, and layers. We then mapped these features to novel text editing capabilities, favouring fast interactions for the most frequent text editing tasks as identified in the taxonomy from Faigley and Witte (1981). Below, we categorize these features based on this taxonomy which differentiates between formal changes (e.g., spelling, tense, number), meaning-preserving changes (e.g., substituting a word with a synonym, reordering words in a sentence), and micro/macro changes. See the accompanying video for a demonstration of the system and its features.
@@ 3.1. Dragging, Resizing, Rotating, and Splitting
Direct interactions are typically reserved for the most frequent operations. In Textoshop, direct interactions map to meaning-preserving changes whose impact is limited to the selection, such as additions, deletions, permutations, distributions, and consolidations (Faigley and Witte, 1981).
@@
[Uncaptioned image] Combining and [Uncaptioned image] Splitting Sentences – Splits a sentence into multiple sentences, or consolidate multiple sentences into one. The interaction is activated through a contextual menu opening after right clicking the text selection.
@@ 3.2. Tools
[Uncaptioned image] Past, [Uncaptioned image] Present, and [Uncaptioned image] Future Tense Changer – Changes the tense of the selection and updates the rest of the text to fix tense agreements. For example, using the future tense changer on “was” in the sentence “Alice was beginning to tire” results in “Alice will begin to tire”.
@@
[Uncaptioned image] Eraser – Removes the selected text and updates the rest of the text accordingly. Erasing makes sure the resulting text remains grammatically correct. Thus it might restructure the sentence and slightly modify other parts. For example, erasing “beginning” in the sentence “Alice was beginning to get very tired” results in “Alice was getting very tired”.
@@ 3.3. Tone Picker
Textoshop adapts traditional colour pickers to allow navigating the space of tones and support meaning-preserving permutations. Essentially, tones are mapped to colours, allowing traditional colour space representations to visually represent the space of tones. Textoshop implements the RGB colour space. A tone in Textoshop is defined along three dimensions (Figure 3). We choose “formality”, “sentiment”, and “complexity” and map them to red, green, and blue based on word-colour associations (Kim et al., 2020). Blending a mix of colours results in the corresponding tone, for instance, a ”purple” tone means formal (red) and simple (blue), and mostly negative (no green), see Figure 3. These dimensions can be edited to explore other tone spaces, if needed. Further, depending on user preferences and tasks, other colour spaces could be more suited.
@@
Tone Wheel – Changes the tone along all three dimensions through one dragging motion in the colour wheel. Tone sliders are updated accordingly. To help users understand the tone wheel, arrows are overlayed, indicating the direction of the strongest change for each of the three tonality dimensions
Changing the Tone After or Before the Selection – Changing the tone after selecting a text passage immediately changes the tone of that passage. This interaction, often referred to as a “noun-verb” interaction (Raskin, 2000), allows rapidly trying different tone variations for that same passage. Or, the tone can be selected beforehand and then applied to the future text selections using the tone brush (section 3.2). This “verb-noun” interaction is better when the goal is to apply the same tone (i.e. verb) to different passages (i.e. noun).
@@ 3.5. Text Layers
In Textoshop, we propose to rely on the familiar layer metaphor. Text from documents can be organized in layers. These layers are similar to layers from drawing software in that they can be viewed as transparent sheets stacked on top of each other. However, unlike drawing layers, text layers store modifications that are anchored in the text from layers underneath it (Figure 5). The reason is that, unlike images, text is positioned relative to other text (e.g., a sentence is positioned after another sentence, but it makes little sense for it to be placed at specific coordinates). This means that the position of the text in a layer is relative to the text of the preceding layers. A modification to any of the preceding layer might reflow the text of the following layers. And layers can hide text below them.
@@
Adding a Layer – Adds an empty layer at the top of the stack. Then, if this layer is selected, all new text modifications are stored in that layer. For example, text from the preceding layers can be modified or removed, and new text can be added, in which case it will be anchored within the text of the preceding layers.
Hiding a Layer – Hides the modifications stored in a layer. For example, if a layer modified the text, hiding this layer restores the original text. And if a layer added some text, hiding the layer removes the added text.
@@ 3.6. Implementation
The Textoshop interface emulates the look and feel of traditional drawing software, except the main canvas is replaced by a text input field (Figure 1). Under the hood, it uses a large-language model as well as more traditional natural language processing techniques. A live demo of the tool is available in supplementary materials [^3] The live demo will be hosted online.
Textoshop is implemented in TypeScript using React (React, 2013) for the interface, Slate.js (SlateJS, 2016) for the text editor, NextUI (NextUI, 2021) for the graphical components, and the OpenAI library (OpenAI, 2023) to execute prompts using OpenAI’s “gpt-4o” model. The source code is available at: https:// [^4] Will be hosted on GitHub..
@@ 3.6.1. Highlighting Text Changes
Due to the difficulties users might have in identifying text changes, we added an animation to highlight character additions and deletions. The animation follows the guidelines from the transitioning technique between text revisions proposed by Chevalier et al. (2010): deleted characters are shown first, in red, and progressively shrink and fade out. Added characters are shown second, in green, and progressively grow and fade in until matching the size of the rest of the text. The whole animation lasts about one second.

### v2

@@ Abstract.
writing, interface metaphors, drawing interaction, LLM, AI
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713862isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Graphical user interfaces
@@
Screenshot of the Textoshop interface with call-outs to show each feature independently. Tools are demonstrated by showing the effect of using the tone brush and smudger. Boolean operations are demonstrated by dragging two text passages. Direct manipulation is demonstrated by showing the effect of resizing and rotating a selection. Text layers are demonstrated by hiding a layer. The tone picker is demonstrated by moving the thumb in the colour wheel.
@@ 1. Introduction
The processes involved in writing share many similarities with those involved in image editing: words form sentences to tell a story, much like pixels come together to create regions that form a complete picture. Just as writers plan and refine a story by outlining the narrative and rearranging the story flow (Flower and Hayes, 1977), visual artists plan and refine a graphical composition by laying out visual elements on a canvas, rearranging them, sometimes organizing them in layers  (Tsandilas et al., 2015). Just as writers combine arguments into sentences and paragraphs (Oshima and Hogue, 2005), visual artists combine shapes into complex drawings (Wood, 2022). And just as writers revise the language and iterate on the wording and structure (Faigley and Witte, 1981), visual artists revise the aesthetics and iterate on the colours and textures (Tsandilas et al., 2015).
@@
To better support writers [^1] By “writers”, we do not necessarily mean professional writers but any individual who engages in writing activities, no matter their expertise, we propose to adapt solutions used in drawing software to help with recurring writing tasks. Specifically, we aim to (1) capitalize on solutions that have been proven to work well and (2) offer solutions most users are already familiar with. Loosely, we view words as pixels, sentences as regions, and tones as colours. This analogy resulted in the design of Textoshop (Figure 1), a text editing software that supports meaning-preserving changes through direct manipulation to move, resize and restructure text selections; a layer management system to organize, display, and edit text in independent views; a sentence composition feature to construct sentences or paragraphs by combining, intersecting, or subtracting ideas or phrases; a tone picker to explore different tones when editing text; and tools including an “eraser” to intelligently remove part of the text, a “tone eyedropper” to copy the tone of a passage, and a “smudge” tool to paraphrase and smoothen transitions between sentences.
@@ 2. Background and Related Work
Traditional text editors such as Microsoft Word offer limited support for editing the text. We first review the issues this creates before discussing solutions proposed in the literature. We then detail the rationale for taking inspiration from drawing software [^2] We use the term “drawing software” to refer to both raster (e.g., Adobe Photoshop) and vector (e.g., Inkscape and Adobe Illustrator) graphics editors. .
@@ 2.1. Issues with Traditional Ways of Editing Text
Second, exploring the space of possible text variations is difficult. A sentence can be rewritten in many ways by varying tone, word order, or length. In fact, rewriting is often considered “the essence of writing” (Zinsser, 2016). Yet, most text editors offer little support to help revise language other than a thesaurus to find synonyms. Some experimental tools leverage LLMs to revise text, but it remains challenging for writers to express their intent with fine granularity (Kim et al., 2024).
@@ 2.2.1. Writing Using Non-linear Text Editors
As early as the development of the World Wide Web, Bolter and Joyce (1987) proposed a writing tool supporting interactive fiction writing using a structural editor with a diagrammatic view. This spurred research into the development of hypertext authoring tools (Kitromili et al., 2019) with which authors write part of the story in an order that might differ from how readers will read it. While these tools were primarily designed to create interactive stories, the idea of being able to write in any order inspired tools to author more traditional and static text. For example, the commercial software Scrivener (Scrivener, 2005) allows writers to compose text in any order by relying on different text areas that can be reorganized to rearrange the final story.
Textoshop also supports non-linear writing: text can be either positioned following lines on a page or freely arranged in 2D in and off the page. The margins can store mood boards, alternatives, and outlines akin to how digital artists leverage the space outside the canvas to preserve elements for reference (Holinaty et al., 2021). Additionally, Textoshop supports layers of text that can be edited independently and combined at will to create different composite stories.
@@ 2.2.2. Revising and Exploring Text Variations
Our work differs in that the space of possible variations is explored through direct manipulation. If a sentence is close to satisfactory, users can explore variations of different lengths or word orders through the resizing and rotating features. They can also move a slider in a two-dimensional space to explore tone nuances, much like visual artists explore colour nuances in a colour picker.
@@ 2.2.3. Reordering and Combining Text Fragments
New writing tools increasingly “shift the writer’s role towards editorial decisions” and manipulating text fragments (Buschek, 2024). For example, Langsmith (Ito et al., 2020) can suggest full sentences from incomplete and rough sentence fragments. Similarly, Rambler (Lin et al., 2024) allows refining and combining dictated train-of-thoughts into full drafts. Other work proposed to manipulate summary representations such as paragraph summaries (Dang et al., 2022) and visual outlines (Zhang et al., 2023) to revise the flow of ideas.
We go beyond simple merging of text passages by exploring novel ways to combine text inspired by boolean operations on shapes, such as subtracting to remove the information in one passage from another, intersecting to preserve only shared information between two passages, and excluding to remove redundant information.
@@ 2.3. Designing Interactions Based on Analogies
Designing software interactions based on such analogies (also referred to as interface metaphors (Neale and Carroll, 1997)) can lead to easily learned and expressive interactions (Rogers et al., 2023, ch 2.4). For example, the direct manipulation interaction paradigm is all about manipulating elements in software as if they were physical objects that can be moved, resized, or distorted (Shneiderman, 1983; Hutchins et al., 1985). Instrumental interaction extends direct manipulation by noting that humans leverage tools when the finger or the hand is not enough (Beaudouin-Lafon, 2000; Beaudouin-Lafon and Mackay, 2000). In the software domain, these tools are interaction instruments that operate on domain objects.
Drawing software combine a canvas analogy with ideas from direct and instrumental interaction. For example, Adobe Photoshop allows elements to be moved, removed, resized, and rotated through rapid and reversible physical actions that mimic how one interacts with physical objects in the real world, just like advocated by direct manipulation. And interaction instruments simulate paint brushes, pencils, or erasers. Other interaction instruments include the colour picker, which is analogous to a paint palette, and layers, which are analogous to transparent sheets (Belge et al., 1993; Andrews, 2006) and reify (Beaudouin-Lafon and Mackay, 2000) the concept of independent manipulation.
The success of drawing software has been inspirational for other software. For example, it is common to find the concept of brushes, colour pickers, and layers adopted by software operating in different domains, such as presentation and video editing. And some of the digital drawing concepts have been adapted successfully, such as using a paintbrush to control the fortune of characters in a story (Chung et al., 2022), mixing prompts just like colours can be mixed to convey visual styles difficult to express with words (Chung and Adar, 2023), or repurposing raster image tools to manipulate colour histograms (Chevalier et al., 2012). Beyond offering more expressive interactions, these analogies help people reason about unfamiliar software by leveraging their accumulated interaction knowledge about digital tools (Renom et al., 2023).
In contrast, text editors evolved from command-line word processors (Mahlow, 2023; Haigh, 2006) and use the analogy of the typewriter and sheet of paper (Kirschenbaum, 2016). In this work, we explore how interaction inspired by the graphical domain could open up new possibilities for writing.
@@ 3. Textoshop
We designed Textoshop by reviewing features found in most raster or vector image editing software and that computer users are likely familiar with, such as colour pickers, drawing tools, and layers. We then mapped these features to novel text editing capabilities, favouring fast interactions for frequent text editing tasks as identified in the taxonomy from Faigley and Witte (1981). Below, we categorize these features based on this taxonomy, which differentiates between formal changes (e.g., spelling, tense, number), meaning-preserving changes (e.g., substituting a word with a synonym, reordering words in a sentence), and micro/macro changes.
@@ 3.1. Dragging, Resizing, Rotating, and Splitting
Direct interactions are typically reserved for frequent operations. In Textoshop, we map these interactions to meaning-preserving changes whose impact is limited to the selection, such as additions, deletions, permutations, distributions, and consolidations (Faigley and Witte, 1981).
@@
Screenshots of the different direct manipulation interactions. Dragging moves the text selection. Resizing is demonstrated in the sentence ”So she was considering in her own mind,” which becomes ”She was pondering”. Rotating is demonstrated in the same sentence and results in ”So in her own mind, she was considering”.
@@
[Uncaptioned image] Combining and [Uncaptioned image] Splitting Sentences – Splits a sentence into multiple, or consolidates multiple sentences into one. The action is in a contextual menu opened by right-clicking a text selection.
@@ 3.2. Tools
[Uncaptioned image] Past, [Uncaptioned image] Present, and [Uncaptioned image] Future Tense Changer – Changes the tense of the selection and updates the rest of the text to fix tense agreements. For example, applying the future tense to “was” in “Alice was beginning to tire” results in “Alice will begin to tire”.
@@
[Uncaptioned image] Eraser – Removes the selected text and updates the rest of the text accordingly. Erasing makes sure the resulting text remains grammatically correct. Thus, it might restructure the sentence and slightly modify other parts. For example, erasing “beginning” in the sentence “Alice was beginning to get very tired” results in “Alice was getting very tired”.
@@ 3.3. Tone Picker
Textoshop adapts traditional colour pickers to allow navigating the space of tones and support meaning-preserving permutations. Essentially, tones are mapped to colours, allowing traditional colour space representations to visually represent the space of tones. Textoshop implements the RGB colour space. A tone in Textoshop is defined along three dimensions (Figure 3). We choose “formality”, “sentiment”, and “complexity” and map them to red, green, and blue based on word-colour associations (Kim et al., 2020). Blending a mix of colours results in the corresponding tone, for instance, a ”purple” tone means formal (red) and simple (blue), and mostly negative (no green), see Figure 3. These tones were selected as defaults because of their orthogonality and their use in previous work on manipulating tones (Jin et al., 2021; Okoso et al., 2024; Huang et al., 2024). These dimensions can be edited to explore other tone spaces, if needed. Further, we note that depending on user preferences and tasks, other colour spaces could be more suited.
@@
Screenshot of the tone picker with arrows indicating what the text would become had the thumb been moved to that position in the space.
@@
Tone Wheel – Changes the tone along all three dimensions through one dragging motion in the colour wheel. Tone sliders are updated accordingly. To help users understand the tone wheel, arrows are overlayed, indicating the direction of the strongest change for each of the three tonality dimensions.
Exploring Other Tones – Tones in the colour picker are input text boxes that can be edited. Users can change the tones to explore new tone spaces, such as “humour” or “Shakespearean”, or extend a scale to have finer granularity, such as going from “formal” to “very formal” to offer 10 levels of formality.
Changing the Tone After or Before the Selection – The tone picker supports both noun-verb and verb-noun command constructions depending on the order of actions (Raskin, 2000, ch 3-3). In a noun-verb construction, the text is selected first (noun), and following changes to the tone (verb) immediately update the passage, to rapidly explore different tone variations. In a verb-noun construction, the tone is selected beforehand (verb) and applied to future text selections (nouns) using the tone brush, to apply the same tone rapidly.
@@ 3.4. Boolean Operations on Sentences
Screenshots with an abstract representation of what boolean operations do. The example tries to merge ”The dog is playing with a bid in the yard” with ”The cat is hunting a bid in the yard” which results in ”The dog is playing with the bird that the cat is hunting in the yard” when uniting, ”The dog is playing” when subtracting, ”The bird is in the yard” when intersecting, and ”the dog is playing while the cat is hunting” when excluding.
@@ 3.5. Text Layers
In Textoshop, we propose to rely on the familiar layer metaphor. Text from documents can be organized in layers. These layers are similar to layers from drawing software in that they can be viewed as transparent sheets stacked on top of each other. However, unlike drawing layers, text layers store modifications that are anchored in the text from layers underneath it (Figure 5). The reason is that, unlike images, text is positioned relative to other text (e.g., a sentence is positioned after another sentence, but it makes little sense for it to be placed at specific coordinates). This means that the position of the text in a layer is relative to the text of the preceding layers. A modification to any of the preceding layers might reflow the text of the following layers. And layers can hide the text below them.
@@
Illustration of the layers. The layers are demonstrated through a 3D visualization of the stack of transparent layers. Depending on which layer is visible, the text changes.
[Uncaptioned image] Adding a Layer – Adds an empty layer at the top of the stack. Then, if this layer is selected, all new text modifications are stored in that layer. For example, text from the preceding layers can be modified or removed, and new text can be added, in which case it will be anchored within the text of the preceding layers.
[Uncaptioned image] Hiding and [Uncaptioned image] Showing a Layer – Hides/shows the modifications stored in a layer. For example, if a layer modified the text, hiding this layer restores the original text. And if a layer added some text, hiding the layer removes the added text.
@@ 3.6. Implementation
The Textoshop interface emulates the look and feel of traditional drawing software, except the main canvas is replaced by a text input field (Figure 1). Under the hood, it uses a large language model as well as more traditional natural language processing techniques.
Textoshop is implemented in TypeScript using React (React, 2013) for the interface, Slate.js (SlateJS, 2016) for the text editor, NextUI (NextUI, 2021) for the graphical components, and the OpenAI library (OpenAI, 2023) to execute prompts using OpenAI’s “gpt-4o” model. A live demo and the source code are available [^3] Live demo and code: https://github.com/m-damien/Textoshop.
@@ 3.6.1. Highlighting Text Changes
To help identify text changes, we added an animation to highlight character additions and deletions. The animation follows the guidelines from the transitioning technique between text revisions proposed by Chevalier et al. (2010): deleted characters appear first, in red, and progressively shrink and fade out. Added characters appear second, in green, and progressively grow and fade in until matching the size of the rest of the text. The whole animation lasts about one second.
3.6.2. Preserving Semantic Meaning
Certain modifications, such as resizing or relocating text, might result in text that is grammatically correct but with a different semantical meaning. We implemented several mechanisms to mitigate the loss of semantic meaning. For example, when resizing, the original sentence is saved and always used for subsequent prompts sent to the model. This means that even if the sentence is reduced to a couple of words, expanding it again will be based on the original longer sentence. Similarly, when relocating a passage, the model is prompted with the sentences surrounding the final location so that the sentence is adjusted to blend into its new location.